the cost of 6 gallons of gasoline is $7.74. how much would 11 gallons of the same gasoline cost?
Answer by checkley71(8403)

6/7.74=11/X OR 6X=(7.74)(11) OR 6X=85.14 OR X=85.14/6 OR X=$14.19